How do you prove that 500=w(w plus 5) when you know that w is 20
Answer by stanbon(25959)

How do you prove that 500=w(w plus 5) when you know that w is 20
-------------
You have to start from what you know and build to where you want to gl.
-------------
w = 20
w+5 = 25
---------------
Multiply the two equations to get:
w(w+5) = 20*25
w(w+5) = 500
===================================
Cheers,
Stan H.